OpenStudy (anonymous):

Sketch the region enclosed by the given curves. Decide whether to integrate with respect to x or y. Draw a typical approximating rectangle and label its height and width. y=7x^2 y^2 = (1/7)x find area of the region

OpenStudy (amistre64):

this sounds familiar ;)

OpenStudy (anonymous):

helloo and yes D: i feel like i did all the steps right, but its not the right answer >< i set both the equations to eachother to find the boundaries, thats the first step right?

OpenStudy (anonymous):

7x^2 = sqrt of (1/7)x i square both sides to get rid of the radical so im left with 7x^4 = (1/7)x i multiply both sides by 7, i get 49x^4 = x subtract x from both sides, you get 49x^4 - x = 0

OpenStudy (amistre64):

can we double int it?

OpenStudy (anonymous):

i believe so

OpenStudy (amistre64):

y = sqrt(1/7x) y=7x^2 {SS} dy dx ; x=[0,1]; y = [7x^2,sqrt(1/7x)] right?

OpenStudy (amistre64):

{S} sqrt(1/7x) - 7x^2 dx ; [0,1]

OpenStudy (anonymous):

how did you get 0 and 1 ?

OpenStudy (amistre64):

sqrt(1/7) 2x sqrt(x/3) - 7/x^3/3

OpenStudy (amistre64):

thats where your graphs intersect at

OpenStudy (anonymous):

to find the intersection, you set the equations equal to eachother right?

OpenStudy (amistre64):

2sqrt(1/21)-7/3 i believe is the area

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

did i miss that up lol

OpenStudy (anonymous):

thats what i did at first, but i got x = 0 and x = cube root of 1/49

OpenStudy (amistre64):

y= 7x^2; and x=7y^2

OpenStudy (amistre64):

i could be wrong, trying tis in my head lol

OpenStudy (anonymous):

lol youre right, but wouldnt we have to set that second equation that y is on one side by itself, not x ?

OpenStudy (amistre64):

y=sqrt(1/7x) ; y = 7x^2 1/7x = 49x^4 49 x^4 - 1/7x = 0 x(49x^3 - 1/7) = 0 x=0 to.....my brain shut off lol

OpenStudy (anonymous):

1/7x(343x^3 - 1) = 0

OpenStudy (anonymous):

x = 0 and x = cube root of 1/343 ?

OpenStudy (amistre64):

wolfram tells me that (1,1/7) is the solution to these; lets try that out :)

OpenStudy (amistre64):

sqrt(1/7*1) = sqrt(1/7) 7*1 = 7....... ack!!

OpenStudy (anonymous):

lol xD

OpenStudy (anonymous):

hmm

OpenStudy (amistre64):

deep breath....... ok

OpenStudy (amistre64):

y=7x^2 and y^2 = 1/7x right?

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

we know zero is a good one

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

y =1 1 7 ---- = 1/7; at x=1/7, y = 1/7 ; thats y=x^2 7 7

OpenStudy (amistre64):

at x = 1/7; we get y = sqrt(1/7 1/7) = sqrt(1/49) = 1/7

OpenStudy (amistre64):

so lets int this thing from 0 to 1/7 :)

OpenStudy (anonymous):

im sorry im a little lost, how did you get 1/7 for x ?

OpenStudy (amistre64):

[S] 7x^2 - sqrt(1/7x) dx ; [0,1/7] ....i guessed and was right :)

OpenStudy (anonymous):

xD awesome haha

OpenStudy (amistre64):

i had an inclination, but couldnt verify it till i just plugged it in ;)

OpenStudy (amistre64):

i know that this is 2 parabolas at right angles to each other; so they only meet at 2 point, to that one works and has to be it lol

OpenStudy (amistre64):

7x^3/3 - 2sqrt((7x)^3)/3 ??

OpenStudy (anonymous):

is the 2aqrt((7x)^3)/3 the same as 2x^(3/2) divided by 21 ? cuz i put sqrt of 1/7 as (1/7)^(1/2) and found the antiderivative like that

OpenStudy (amistre64):

sqrt(1/7)* sqrt(x) = sqrt(1/7)* 2sqrt(x^3)/3 7x^3/3 - 2sqrt(1/7x^3)/3 right?

OpenStudy (amistre64):

the sqrt(1/7) is actually a constant, so we can pull it out of the way and just do sqrt(x)

OpenStudy (amistre64):

sqrt(ab) = sqrt(a) * sqrt(b)

OpenStudy (anonymous):

ohhh i see

OpenStudy (amistre64):

1/49(3) - 2sqrt(1/7^4)/3 should be the area

OpenStudy (anonymous):

how comes its (1/7^4) ?

OpenStudy (amistre64):

1 2sqrt(1/2401) --- - ------------ 147 3 (1/7) * (1/7)^3 = 1/7^4

OpenStudy (amistre64):

as long as i integrated it right, and im pretty sure i did :)

OpenStudy (amistre64):

dbl check it tho ;)

OpenStudy (anonymous):

im going over it right now :D

OpenStudy (amistre64):

\[\int\limits_{} (7x^2 - \sqrt{\frac{1}{7}x} )dx\]

OpenStudy (amistre64):

id take the absolute value of that :)

OpenStudy (anonymous):

thats what i got too :) and i jus ttake the absolute value of that since its negative right ?

OpenStudy (amistre64):

correct; just means we got the wrong order

OpenStudy (anonymous):

gotcha. thank you once again. always coming to my rescue :D haha youre the best !